OpenStudy (anonymous):

I got 6 dollers am i right Determine which estimate is most reasonable. Francesca purchases a box of cereal for $2.98, hamburger for $4.12, tomatoes for $3.07, and milk for $1.96. She has $17 with her. About how much change should she expect to receive?

OpenStudy (phi):

cereal for $2.98, round to 3 hamburger for $4.12, 4 tomatoes for $3.07, 3 and milk for $1.96. 2 3+4+3+2= 7+5=12 About how much change should she expect?

OpenStudy (anonymous):

12?

OpenStudy (phi):

12 is how much she spends.

OpenStudy (phi):

they want to know how much money she has left if she started with 17 and spends 12

OpenStudy (anonymous):

5?

OpenStudy (phi):

yes